# Non-stationary Gaussian random fields on hypersurfaces: Sampling and strong error analysis

### v1

@@ B.2. Norm and error estimates of discrete operators
B.3. Proof of Equation 24
@@
$\mathcal{E}_{h}=\big{\|}\frac{1}{2\pi i}\int_{\Gamma}\gamma(z)\mathcal{F}(z)\,\mathrm{d}z\big{\|}_{L^{2}(\mathcal{M}_{h})},$
@@
(64) $\mathcal{E}_{h}$ $\leq\frac{1}{2\pi}\int_{\Omega_{+}}|\gamma(g_{+}(t))|\left\|\mathcal{F}(g_{+}(t))\right\|_{L^{2}(\mathcal{M}_{h})}\,\mathrm{d}t+\frac{\delta_{0}}{2\pi}\int_{\Omega_{0}}|\gamma(g_{0}(t))|\|\mathcal{F}(g_{0}(t)\|_{L^{2}(\mathcal{M}_{h})}\,\mathrm{d}t$
$\qquad+\frac{1}{2\pi}\int_{\Omega_{-}}|\gamma(g_{-}(t))|\left\|\mathcal{F}(g_{-}(t))\right\|_{L^{2}(\mathcal{M}_{h})}\,\mathrm{d}t,$
@@
We first bound $\left\|\mathcal{F}_{1}(z)\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}$. Using successively Equation 23 and the geometric estimates in Bonito et al. (2018, Corollary 2.2) results in
@@
To bound $\big{\|}\mathcal{F}_{2}(z)\tilde{f}\big{\|}_{L^{2}(\mathcal{M}_{h})}$, we apply Equations 53 and 55 to obtain
@@
and with Equation 52 for any $p\in[-1,1]$ ($p=-q$)
@@
Using Equations 67 and 66 with $p=\min\{\alpha+d/4;1\}$ together with Equation 65 gives
@@
which yields in turn (since $\gamma$ is an $\alpha$-power spectral density)

### v2

@@ B.2. Norm estimates
Finally, we recall the Bramble-–Hilbert lemma (cf. [6, Equation (4.8)]), which is used in several proofs in this paper.
Lemma B.4 (Bramble-–Hilbert lemma).
For any $t\in[0,2]$ and any $\varphi\in H^{t}(\mathcal{M})$,
(69) $\|(I-P_{h})\varphi\|_{L^{2}(\mathcal{M})}\lesssim h^{t}\|\varphi\|_{H^{t}(\mathcal{M})}\lesssim h^{t}\|\mathcal{L}^{t/2}\varphi\|_{L^{2}(\mathcal{M})},$
where we used the equivalence of Sobolev norms dot-spaces norms in the last inequality.
B.3. Proof of Lemma 2.5
Proof.
Let $h\in(0,h_{0})$, $z\in\Gamma$. Let $f\in L^{2}(\mathcal{M})$ and let $p\in[0,1]$ such that $\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}<\infty$. Finally, let $\beta\in[0,1]$ such that $p\in[0,(1+\beta)/2]$. We then define $\mathcal{F}_{h}(z)$ by $\mathcal{F}_{h}(z)=(z-\mathcal{L}_{h})^{-1}P_{h}-P_{h}(z-\mathcal{L})^{-1}$. Note then
$\mathcal{F}_{h}(z)$ $=(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}(\mathcal{L}_{h}^{-1}P_{h}(z-\mathcal{L})\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}(z-\mathcal{L}_{h})P_{h}\mathcal{L}^{-1})\mathcal{L}(z-\mathcal{L})^{-1}$
$=(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}(P_{h}\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\mathcal{L}(z-\mathcal{L})^{-1}$
Hence,
$\|\mathcal{F}_{h}(z)f\|_{L^{2}(\mathcal{M})}$ $=\|(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}(P_{h}\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\mathcal{L}(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}$
$=\|\mathcal{L}_{h}^{1-(1-\beta)/2}(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}^{(1-\beta)/2}(P_{h}\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\mathcal{L}(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}$
$\lesssim|z|^{-(1-\beta)/2}\|\mathcal{L}_{h}^{(1-\beta)/2}(P_{h}\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\mathcal{L}(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}$
where we used Equation 58 with $s=(1-\beta)/2\in[0,1/2]$ to derive the inequality. Note then that, using Equation 59, we have
$\|\mathcal{F}_{h}(z)f\|_{L^{2}(\mathcal{M})}$ $\lesssim|z|^{-(1-\beta)/2}h^{2\beta}\|\mathcal{L}^{-(1-\beta)/2}\mathcal{L}(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}$
$=h^{2\beta}|z|^{-(1-\beta)/2}\|\mathcal{L}^{(1+\beta)/2}(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}$
$=h^{2\beta}|z|^{-(1-\beta)/2}\|\mathcal{L}^{(1+\beta)/2-p}(z-\mathcal{L})^{-1}\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}$
Using then Equation 57 with $s=(1+\beta)/2-p\in[0,1]$, we retrieve
$\|\mathcal{F}_{h}(z)f\|_{L^{2}(\mathcal{M})}$ $\lesssim h^{2\beta}|z|^{-(1-\beta)/2}|z|^{-(1-(1+\beta)/2+p)}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}$
$=h^{2\beta}|z|^{-(1-\beta+p)}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}.$
This concludes the proof. ∎
B.4. Proof of Equation 26
@@
$\mathcal{E}_{h}=\big{\|}\frac{1}{2\pi i}\int_{\Gamma}\gamma(z)\mathcal{F}(z)\tilde{f}\,\mathrm{d}z\big{\|}_{L^{2}(\mathcal{M}_{h})},$
@@
(70) $\mathcal{E}_{h}$ $\leq\frac{1}{2\pi}\int_{\Omega_{+}}|\gamma(g_{+}(t))|\left\|\mathcal{F}(g_{+}(t))\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}\,\mathrm{d}t+\frac{\delta_{0}}{2\pi}\int_{\Omega_{0}}|\gamma(g_{0}(t))|\|\mathcal{F}(g_{0}(t))\tilde{f}\|_{L^{2}(\mathcal{M}_{h})}\,\mathrm{d}t$
$\qquad+\frac{1}{2\pi}\int_{\Omega_{-}}|\gamma(g_{-}(t))|\left\|\mathcal{F}(g_{-}(t))\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}\,\mathrm{d}t,$
@@
We first bound $\left\|\mathcal{F}_{1}(z)\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}$. Using successively Equation 25 and the geometric estimates in [5, Corollary 2.2] results in
@@
To bound $\big{\|}\mathcal{F}_{2}(z)\tilde{f}\big{\|}_{L^{2}(\mathcal{M}_{h})}$, we apply Equations 61 and 63 to obtain
@@
$=|z|^{-1/2}h^{2}\left\|\mathcal{L}_{h}^{(1+p)/2}(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}^{-p/2}\tilde{f}\right\|_{L^{2}(\mathcal{M})}$
and with Equation 58 (applied with $s=(1+p)/2$)
@@
Using Equations 73 and 72 with $p=\min\{\alpha+d/4;1\}$ together with Equation 71 gives
@@
which yields in turn (since $\gamma$ is an $\alpha$-amplitude spectral density)